The length of a rectangle is 10 in. more than twice its width. If the perimeter of the rectangle is 32 in., find the width of the rectangle.
Answer by BurningFlame_marc(77)

Hello!

First, let x=width 2x+10=length
Equation:
2x+2(2x+10)=32
Solution:
2x+4x+20=32
6x=12
x=2
Final Answer: the width is 2 inches.